Calculate the sensitivity, specificity, positive predictive value, and negative predictive value for the Test Screen.

                                   Medical Assessment

Test Screen              Disease          No Disease Total

Possible Disease   30                    20                    50

No Disease              10                    40                    50

Total                          40                    60                    100
Answer #1

Sensitivity = P ( Test says Dz | Actual Disease) = 30/40 = 0.75

Specificity = P ( Test says No Dz | Actual No Disease) = 40/60 = 0.666

Positive Pred Value = P ( Actual Disease | Test says Dz ) = 30 / 50 = 0.60

Negative Pred Value = P ( Actual No Dz | Test Says No Dz ) = 40/50 = 0.8

Vote if this is useful.